OpenStudy (anonymous):

Find the indicated output for this function. if h(x)=x^3-2, find h(2)

OpenStudy (anonymous):

h(2)=(2)^3 -2 h(2)=8-2 h(2)=6

OpenStudy (anonymous):

so 6 is the answer right